Calculate the solubility of lead chloride, PbCl2 in units of grams per liter. Ksp(PbCl2) = 1.7×10-5.

solubility =  g/L
Answer #1

Let the solubility of pbcl2 is X M

PbCl2. = Pb2+ + 2cl-

X X. 2X

Ksp = X (2X)2= 4X3

4X3= 1.7 E-5

X3= 1.7/4 E-5

X= 1.619E-2

Solubility in g/L= 0.01619× 278.1= 4.50g/L